OpenStudy (anonymous):

um.....how do u multiply fraction.....<3

Parth (parthkohli):

You multiply numerators and denominators \(\heartsuit \)

Parth (parthkohli):

\( \color{Black}{\Rightarrow \Large {a \over b} \times {c \over d} = {ac \over bd} }\)

Parth (parthkohli):

Example will work here :P \( \color{Black}{\Rightarrow \Large {1 \over 2} \times {1 \over 2} = {1 \times 1 \over 2 \times 2} = {1 \over 4} }\)

OpenStudy (anonymous):

thxs